OpenStudy (anonymous):

What is the equation of the line passing through (–6, 1) and perpendicular to the line y = –3x + 1 in slope-intercept form?

OpenStudy (anonymous):

The slope of a perpendicular line is always the negative reciprocal of the line you are given. So you know that the slope of your line is going to be 1/3, which is -3 flipped and the sign reversed. y=1/3x+b Plug in the points you are given to find your y intercept (b): 1=1/3(-6)+b b=3 So your equation is: y=1/3x+3